An automobile leaves New York for Chicago on a road paralleling a railroad track at a rate averaging 26 miles per hour. Six hours later an express train leaves Chicago, averaging 50 miles per hour. In how many hours will the train overtake the automobile?
Answer by fcabanski(1391)

Distance = rate * time.


The car (who is driving it that slowly?) travels D = 26*6 = 156 miles in six hours.


The train overtakes that car at 50-26 = 24 MPH, and must make up 156 miles. 156 = 24*t and thus 156/24 = t = 6.5 hours.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)